OpenStudy (anonymous):

i need you :) Solve the equation x-4/x-1 - x+3/x =0

OpenStudy (anonymous):

Do you know how to add the fractions together?

OpenStudy (anonymous):

not :D

OpenStudy (uri):

\[x-4/ x-1 - x+3/x =0\] Take L.c.m and L.c.m will be x (x-1)

OpenStudy (anonymous):

Ok.\[\frac{x-4}{x-1}-\frac{x+3}{x}=0\]In order to add the fractions together, multiply the first fraction by the denominator of the second fraction over itself, and multiply the second fraction by the denominator of the first fraction over itself. So\[\frac{x-4}{x-1}\times \frac{x}{x}-\frac{x+3}{x}\times \frac{x-1}{x-1}=0\]\[\frac{x^{2}-4x}{x(x-1)}-\frac{x^{2}+3x}{x(x-1)}=0\]Now you simply add the numerators together.\[\frac{2x^{2}-x}{x(x-1)}=0\]Now in order to solve the equation you only need to solve\[x(2x-1)=0\]making sure that you exclude any solution that is also a solution of \[x(x-1)=0\]Does that make sense?

OpenStudy (anonymous):

thank you :)